# BondMatcher: H-Bond Stability Analysis in Molecular Systems

### v1

@@ 2.3 Molecular vibrations
Molecular vibrations are intrinsic behaviors of molecular systems. They describe the relative motion of atoms (occurring at all temperatures) which does not change the position of the molecular center of mass. A system composed of $n_{a}$ atoms is a subject to $n_{m}=3n_{a}-k$ vibrational degrees of freedom, with $k=3+3$ in non-linear molecules (to discard the 3 translational and 3 rotational degrees of freedom of the molecular center of mass). In particular, for a water hexamer, (H2O)6, $n_{a}=18$ and $n_{m}=48$. The complex motion induced by these degrees of freedom can be decomposed into simpler components, referred to as the normal modes of vibrations. Each mode describes the motion of all atoms occurring with a specific frequency and phase.
Specifically, the equilibrium structure of each isomer, noted ${\mathcal{R}}^{0}$, was optimized with the PBE0/TZ2P model (Sec. 2.2) and the harmonic vibrational modes, noted ${\mathcal{Q}}$, were computed via normal mode analysis (NMA), a standard method in chemistry for studying molecular vibrations under the harmonic oscillator approximation [95]. Technically, the displacement of a nucleus $i$ along each of the $n_{m}$ vibrational normal vectors ${\mathcal{Q}}_{j}\in\mathbb{R}^{3n_{a}}$ is given by:
@@
Next, let $\alpha_{k}$ be an integer in the range $(0,20)$. Then, the position of a nucleus $i$ in the perturbed geometry $k$ of the vibration mode $j$ is:
@@
where $\mathcal{A}_{h}$ is a classical estimation of the turning point of the harmonic oscillatory motion [30].
@@ 3.1 Chemical features of interest specification
Given the electron density $\rho:\mathcal{K}\rightarrow\mathbb{R}$ of a water hexamer, following QTAIM (detailed in Sec. 1.2), the goal of our analysis is to extract a geometrical representation of the following features of interest: oxygen atoms, hydrogen atoms, covalent bonds, H-bonds.
@@ 3.2 Topological characterization
Once the above two pre-processes are completed (to remove spurious minima and $1$-saddles), the extremum graph is computed by collecting the unstable sets (Sec. 1.3) of each remaining $1$-saddle. Formally, the extremum graph of $\rho^{\prime}$, noted $\mathcal{E}(\rho^{\prime})$, is a graph with a set of nodes $\mathcal{N}(\rho^{\prime})$ and a set of arcs $\mathcal{A}(\rho^{\prime})$, such that each node $n\in\mathcal{N}(\rho^{\prime})$ represents exactly one minimum of $\rho^{\prime}$, and each arc $a\in\mathcal{A}(\rho^{\prime})$ represents the unstable set of a $1$-saddle of $\rho^{\prime}$ linking two of its minima.
@@ 3.3 From topological to chemical features
Since oxygen and hydrogen atoms have notably different atomic masses, these can be easily distinguished based on their opposite electron density values ($\rho^{\prime}$), typically in our database around $-8$ for oxygen atoms and above $-1$ for hydrogen atoms. Thus, we classify all atoms with an opposite electron density value below (respectively above) $-4$ as oxygen atoms (respectively hydrogen atoms).
Similarly, since they involve strong electrostatic interactions, covalent bonds can be easily distinguished from non-covalent bonds based on the opposite electron density value ($\rho^{\prime}$) of their corresponding bond critical point. Specifically, we classify as covalent bond (respectively non-covalent bond) an unstable set from the extremum graph which connects an oxygen atom to a hydrogen atom, and whose bond critical point has an opposite electron density value below (respectively above) $-0.1$. Note that this isovalue of opposite electron density ($-0.1$) is often selected for representing the geometry of a molecule via isosurfacing (Fig. 3). In certain cases, unstable sets may connect two oxygen atoms, which does not characterize a priori any chemically relevant interaction in this context. We will refer to such connections as misconnected H-bonds (see Sec. 5.2).
@@
Figure 5: Geometry-aware partial isomorphism estimation between the extremum graphs of the Prism water hexamer at equilibrium (top) and under large vibrational displacement (bottom). (a) First, an optimal assignment matches the nodes together (colors), in order to minimize the sum of their 3D mismatches (Eq. 1). (b) This matching is extended to the arcs (colors) by matching arcs together if their nodes are matched by the optimal node assignment, possibly yielding a partial isomorphism (unmatched arcs are shown in white). (c) For this ensemble (only two graphs), the bond occurrence rate reports a value of $1$ (red) for all arcs involved in the partial isomorphism and of $0.5$ (black) for the others.
Figure 6: Bond occurrence analysis for the Prism (vibration mode 00). (a) Extremum graphs for the input vibrational displacements (black to red: negative to positive translations, equilibrium: white). (b) Partial isomorphisms to the equilibrium state (colored arcs, white: unmatched arcs). (c) Bond occurrence rate (black to red). (d-h) Bond occurrence rate for various configurations, as a function of the vibrational displacement: (d) missing H-bond ($\alpha_{0}$), (e) misconnected H-bonds ($[\alpha_{1},\alpha_{2}]$), (f) valid H-bonds ($[\alpha_{3},\alpha_{12}]$), (g) misconnected H-bonds ($[\alpha_{13},\alpha_{17}]$), (h) missing H-bonds ($[\alpha_{18},\alpha_{20}]$). Valid H-bonds occur around the equilibrium, while misconnected and missing H-bonds respectively occur for intermediate and large displacements.
@@ 4 Analysis of Ensembles of Extremum Graphs
This section describes our approach for estimating a geometry-aware partial isomorphism between two extremum graphs, representing two distinct electron densities. This partial isomorphism is then exploited in the definition of our stability measure (called bond occurrence rate).
@@ 4.1 Overview
The goal of the approach documented in this section is to establish a reliable correspondence between these two extremum graphs. For this, we first establish a correspondence $\phi_{\mathcal{N}_{i\rightarrow j}}:\mathcal{N}(\rho^{\prime}_{i})\rightarrow\mathcal{N}(\rho^{\prime}_{j})$ between the nodes of the extremum graphs (see Sec. 4.2 and Fig. 5(a)), prior to extending it to their arcs, in the form of an induced partial isomorphism $\phi_{\mathcal{A}_{i\rightarrow j}}:\mathcal{A}(\rho^{\prime}_{i})\rightarrow\mathcal{A}(\rho^{\prime}_{j})$ (see Sec. 4.3 and Fig. 5(b)). To do so, we specifically exploit the fact that our database models water hexamers. Therefore, at this stage, each extremum graph counts exactly $18$ nodes (representing the $3$ atoms of the $6$ water molecules). Then, we have $|\mathcal{N}(\rho^{\prime}_{i})|=|\mathcal{N}(\rho^{\prime}_{j})|$, which is a valuable simplifying hypothesis.
@@ 4.2 Optimal node assignment
While the original molecule specifications (i.e., the atom locations) are known for both opposite densities $\rho^{\prime}_{i}$ and $\rho^{\prime}_{j}$ (this is the input of the DFT computation, Sec. 2), in practice, the locations of their resulting minima do not exactly coincide to this specification due to both spatial quantization artifacts and the imbalance between the atomic masses of the oxygen and hydrogen atoms (which displaces the minima, especially those associated to hydrogen atoms). Then, a robust algorithm needs to be considered to establish a reliable correspodence between the node sets $\mathcal{N}(\rho^{\prime}_{i})$ and $\mathcal{N}(\rho^{\prime}_{j})$. For this, we consider the following balanced assignment problem (the simplest form of optimal transport [70]).
Let $\phi_{i\rightarrow j}:\mathcal{N}(\rho^{\prime}_{i})\rightarrow\mathcal{N}(\rho^{\prime}_{j})$ be a bijection between the node sets $\mathcal{N}(\rho^{\prime}_{i})$ and $\mathcal{N}(\rho^{\prime}_{j})$. Then, let $E(\phi_{i\rightarrow j})$ be the following energy term:
@@
The term $E(\phi_{i\rightarrow j})$ simply evaluates the sum of the distances in the original 3D space between the minima of $\rho^{\prime}_{i}$ and their images by $\phi_{i\rightarrow j}$ (i.e., the sum of the geometrical mismatches between minima). Then, we select as final correspondence $\phi_{\mathcal{N}_{i\rightarrow j}}$ a global minimizer of $E(\phi_{i\rightarrow j})$:
@@ 4.3 Induced partial isomorphism
The optimal node assignment computed in the previous section is extended into an induced partial isomorphism $\phi_{\mathcal{A}_{i\rightarrow j}}$ as follows. Let $a_{i}\in\mathcal{A}(\rho^{\prime}_{i})$ and $a_{j}\in\mathcal{A}(\rho^{\prime}_{j})$ be two arcs of $\mathcal{E}(\rho^{\prime}_{i})$ and $\mathcal{E}(\rho^{\prime}_{j})$ respectively. Then, we assign $a_{i}$ to $a_{j}$, i.e., $\phi_{\mathcal{A}_{i\rightarrow j}}(a_{i})=a_{j}$, if and only if:
@@
Figure 7: H-bond path configurations for the Prism (vibration mode 00), according to their unstable sets (white: saddles, red: minima, black curves: unstable sets). A 2D slide of the electron density is shown in the background with three isolines. The corresponding partial isomorphisms are shown with colors in the insets. (a) At equilibrium ($\alpha_{10}$), the unstable set of the saddle $A$ connects it to hydrogen (left) and oxygen (right) atoms. (b) Under intermediate displacement ($\alpha_{2}$), the saddle $A$ increases its $\rho^{\prime}$ value (white isoline). The geometry of its unstable set is altered such that it now connects it to two oxygen atoms. We reter to this configuration as a misconnected H-bond. (c) Under large displacement ($\alpha_{0}$), the saddle $A$ disappears. The electron density field has been sufficiently modified by the vibrational displacement to cancel the corresponding saddle-saddle persistence pair [17]. We reter to this configuration as a missing H-bond.
@@ 5 Results
This section presents experimental results obtained with a C++ implementation of our approach, implemented as a module for TTK [91, 9].
5.1 Database specification
The complete specification, with download links and preparation scripts, for the database of electron densities generated as documented in Sec. 2 is available at this address: https://github.com/thom-dani/BondMatcher.
@@ 5.2 Geometrical interpretation
Running the topological data analysis pipeline described in Sec. 3 yields a database of extremum graphs. Fig. 7 presents a classification of the H-bond configurations that we have observed in this database, according to their associated unstable sets (Sec. 1.3). This figure shows that, under vibrational displacement, the bond critical point (Sec. 1.2) associated to an H-bond path gradually increases its $\rho^{\prime}$ value. Depending on the amplitude of the vibrational displacement, this can result in two degenerate configurations. First, a misconnected H-bond, Fig. 7(b), occurs when the $\rho^{\prime}$ value of the bond critical point increased to an intermediate level, such that the geometry of its unstable set is altered to the point that it connects two oxygen atoms together (a configuration which is not chemically relevant in this context). Second, a missing H-bond, Fig. 7(c), occurs when the $\rho^{\prime}$ value of the bond critical point increased so much that the corresponding saddle-saddle pair has been canceled [17]. This figure suggests that the transition from a valid H-bond path at equilibrium to a missing one is gradual (via a misconnection step), which may relate to experimental studies discussing the role of oxygen-oxygen contractions on vibrations of liquid water [97].
Fig. 6 confirms this observation by inspecting the entire set of $21$ vibrational displacements for the fist vibrational mode of the Prism. Specifically, it shows that missing H-bonds occur at the extremities of the displacement spectrum ($\alpha_{0}$ and $[\alpha_{18},\alpha_{20}]$) and that misconnected H-bonds occur within immediately adjacent intervals ($[\alpha_{1},\alpha_{3}]$ and $[\alpha_{13},\alpha_{17}]$). This confirms that the transition from valid (around the equilibrium), to misconnected, to missing H-bond is indeed gradual along the vibration. Fig. 6(c) also shows the bond occurrence rate (Sec. 4.4) for these configurations. It reports lower rates for misconnected H-bonds, indicating the transitional aspect of this configuration.
@@
Fig. 8 investigates the relation between the unstable bonds of the Prism (relative to vibrations) and topological persistence [23], a generic importance measure often used to estimate prominence and stability in Topological Data Analysis [22]. In particular, this figure shows that, in this example, the bonds identified as unstable are defined relatively to saddles which are also involved in the least persistent saddle-saddle pairs (similar observations, not reported here, were made for the Cage isomer). This indicates that such low-persistence saddles are more likely to undergo, under vibrations, the transitions observed in Fig. 7. Note however, that in theory, instability can still be observed with unstable sets defined relative to the most persistent saddles, as illustrated in Fig. 4. Then, although persistence and bond stability appear to be correlated in our experiments, it is not necessarily the case in generality.
@@
Figure 8: Observing the persistence [23] and bond occurrence rate on the Prism (equilibrium). (a) Saddle persistence pairs (persistence: black to red). (b-c) Bond occurrence rate (black to red), for mode 00 (b) and mode $2$ (c). In this example, the saddles defining unstable H-bonds (arrows) are also involved in the least persistent saddle pairs (black, gray, pink).

### v2

@@ 2.3 Molecular vibrations
Molecular vibrations are intrinsic behaviors of molecular systems. They describe the relative motion of atoms (occurring at all temperatures) which does not change the position of the molecular center of mass. In the following, although our molecular vibration modeling follows the typical approach in chemistry [82], we still specify its details in order to precisely document our data generation process, thereby enabling reproducibility. A system composed of $n_{a}$ atoms is a subject to $n_{m}=3n_{a}-k$ vibrational degrees of freedom, with $k=3+3$ in non-linear molecules (to discard the 3 translational and 3 rotational degrees of freedom of the molecular center of mass). In particular, for a water hexamer, (H2O)6, $n_{a}=18$ and $n_{m}=48$. The complex motion induced by these degrees of freedom can be decomposed into simpler components, referred to as the normal modes of vibrations. Each mode describes the motion of all atoms occurring with a specific frequency and phase.
The equilibrium structure of each isomer, noted ${\mathcal{R}}^{0}$, was optimized with the PBE0 TZ2P model (Sec. 2.2) and the harmonic vibrational modes, noted ${\mathcal{Q}}$, were computed via normal mode analysis (NMA), a standard method for studying molecular vibrations under the harmonic oscillator approximation [82]. The displacement of a nucleus $i$ along each of the $n_{m}$ vibrational normal vectors ${\mathcal{Q}}_{j}\in\mathbb{R}^{3n_{a}}$ is then given by:
@@
Next, let $\alpha_{k}$ be an integer in the range $(0,20)$, which constitutes a sufficient sampling for our case study (to enable the observation of clear transitions in H-bond configurations, see Sec. 5.3). The position of a nucleus $i$ in the perturbed geometry $k$ of the vibration mode $j$ is:
@@
where $\mathcal{A}_{h}$ is a classical estimation of the turning point of the harmonic oscillatory motion [26]. Fig. 1 illustrates the resulting vectors on a water molecule at mode $j=0$, for extreme displacements (i.e., $k\in\{0,20\}$).
@@ 3.1 Chemical features of interest specification
Given the electron density $\rho:\mathcal{K}\rightarrow\mathbb{R}$ of a water hexamer, and following QTAIM (detailed in Sec. 1.2), the goal of our analysis is to extract a geometrical representation of the following features of interest: oxygen atoms, hydrogen atoms, covalent bonds, H-bonds.
@@ 3.2 Topological characterization
Once the above two pre-processes are completed (to remove spurious minima and $1$-saddles), the extremum graph [17] is computed by collecting the unstable sets (Sec. 1.3) of each remaining $1$-saddle. Formally, the extremum graph of $\rho^{\prime}$, noted $\mathcal{E}(\rho^{\prime})$, is a graph with a set of nodes $\mathcal{N}(\rho^{\prime})$ and a set of arcs $\mathcal{A}(\rho^{\prime})$, such that each node $n\in\mathcal{N}(\rho^{\prime})$ represents exactly one minimum of $\rho^{\prime}$, and each arc $a\in\mathcal{A}(\rho^{\prime})$ represents the unstable set of a $1$-saddle of $\rho^{\prime}$ linking two of its minima.
@@ 3.3 From topological to chemical features
Since oxygen and hydrogen atoms have notably different atomic masses, these can be easily distinguished based on their opposite electron density values ($\rho^{\prime}$), typically in our database around $-8$ for oxygen atoms and above $-1$ for hydrogen atoms. Thus, we classify as oxygen atoms all the atoms with an opposite electron density value below $-4$, and as hydrogen atoms all the atoms above this value.
Similarly, since they involve strong electrostatic interactions, covalent bonds can be easily distinguished from non-covalent bonds based on the opposite electron density value ($\rho^{\prime}$) of their corresponding bond critical point. Specifically, we classify as covalent bond an unstable set from the extremum graph which connects an oxygen atom to a hydrogen atom, and whose bond critical point has an opposite electron density value below $-0.1$. Symmetrically, unstable sets connecting an oxygen atom to a hydrogen atom, and whose bond critical point has an opposite electron density value above $-0.1$ are classified as non-covalent bonds. Note that this isovalue of opposite electron density ($-0.1$) is often selected for representing the geometry of a molecule via isosurfacing (Fig. 3). In certain cases, unstable sets may connect two oxygen atoms, which does not characterize a priori any chemically relevant interaction in this context. We will refer to such connections as misconnected H-bonds (see Sec. 5.1).
@@
Figure 5: Geometry-aware partial isomorphism estimation between the extremum graphs of the Prism water hexamer at equilibrium (top) and under large vibrational displacement (bottom). (a) First, an optimal assignment matches the nodes together (colors), in order to minimize the sum of their 3D distances (Eq. 1). (b) This matching is extended to the arcs (colors) by matching arcs together if their nodes are matched by the optimal node assignment, possibly yielding a partial isomorphism (unmatched arcs are shown in white). (c) For this ensemble (only two graphs), the bond occurrence rate reports a value of $1$ (red) for all arcs involved in the partial isomorphism and of $0.5$ (black) for the others.
Figure 6: Bond occurrence analysis for the Prism (vibration mode 00). (a) Extremum graphs for the input vibrational displacements (black to red: negative to positive translations, equilibrium: white). (b) Partial isomorphisms to the equilibrium state (colored arcs, white: unmatched arcs). (c) Bond occurrence rate (black to red). (d-h) Bond occurrence rate for various configurations, as a function of the vibrational displacement: (d) missing H-bond ($\alpha_{0}$), (e) misconnected H-bonds ($[\alpha_{1},\alpha_{2}]$), (f) valid H-bonds ($[\alpha_{3},\alpha_{12}]$), (g) misconnected H-bonds ($[\alpha_{13},\alpha_{17}]$), (h) missing H-bonds ($[\alpha_{18},\alpha_{20}]$). Valid H-bonds occur around the equilibrium, while misconnected and missing H-bonds (arrows) occur for intermediate and large displacements.
@@ 4 Analysis of Ensembles of Extremum Graphs
This section describes our approach for estimating a geometry-aware partial isomorphism between two extremum graphs, representing two distinct electron densities. This partial isomorphism is later exploited in the definition of our stability measure (called bond occurrence rate).
@@ 4.1 Overview
The goal of the approach documented in this section is to establish a reliable correspondence between these two extremum graphs. For this, we first establish a correspondence $\phi_{\mathcal{N}_{i\rightarrow j}}:\mathcal{N}(\rho^{\prime}_{i})\rightarrow\mathcal{N}(\rho^{\prime}_{j})$ between the nodes of the extremum graphs (see Sec. 4.2 and Fig. 5(a)), prior to extending it to their arcs, in the form of an induced partial isomorphism $\phi_{\mathcal{A}_{i\rightarrow j}}:\mathcal{A}(\rho^{\prime}_{i})\rightarrow\mathcal{A}(\rho^{\prime}_{j})$ (see Sec. 4.3 and Fig. 5(b)). To do so, we specifically exploit the fact that our database models water hexamers. Therefore, at this stage, each extremum graph counts exactly $18$ nodes (representing the $3$ atoms of the $6$ water molecules) and we have $|\mathcal{N}(\rho^{\prime}_{i})|=|\mathcal{N}(\rho^{\prime}_{j})|$, which is a valuable simplifying hypothesis.
@@ 4.2 Optimal node assignment
While the original molecule specifications (i.e., the atom locations) are known for both opposite densities $\rho^{\prime}_{i}$ and $\rho^{\prime}_{j}$ (this is the input of the DFT computation, Sec. 2), in practice, the locations of their resulting minima do not exactly coincide to this specification due to both spatial quantization artifacts and the imbalance between the atomic masses of the oxygen and hydrogen atoms (which displaces the minima, especially those associated to hydrogen atoms). Thus, a robust algorithm needs to be considered to establish a reliable correspodence between the node sets $\mathcal{N}(\rho^{\prime}_{i})$ and $\mathcal{N}(\rho^{\prime}_{j})$. For this, we consider the following balanced assignment problem (the simplest form of optimal transport [60]).
Let $\phi_{i\rightarrow j}:\mathcal{N}(\rho^{\prime}_{i})\rightarrow\mathcal{N}(\rho^{\prime}_{j})$ be a bijection between the node sets $\mathcal{N}(\rho^{\prime}_{i})$ and $\mathcal{N}(\rho^{\prime}_{j})$. Next, let $E(\phi_{i\rightarrow j})$ be the following energy term:
@@
The term $E(\phi_{i\rightarrow j})$ simply evaluates the sum of the distances in 3D space between the minima of $\rho^{\prime}_{i}$ and their images by $\phi_{i\rightarrow j}$ (intuitively, the sum of the geometrical mismatches between minima). Finally, we select as correspondence $\phi_{\mathcal{N}_{i\rightarrow j}}$ a global minimizer of $E(\phi_{i\rightarrow j})$:
@@ 4.3 Induced partial isomorphism
The optimal node assignment computed in the previous section is extended into an induced partial isomorphism $\phi_{\mathcal{A}_{i\rightarrow j}}$ as follows. Let $a_{i}\in\mathcal{A}(\rho^{\prime}_{i})$ and $a_{j}\in\mathcal{A}(\rho^{\prime}_{j})$ be two arcs of $\mathcal{E}(\rho^{\prime}_{i})$ and $\mathcal{E}(\rho^{\prime}_{j})$ respectively. Next, we assign $a_{i}$ to $a_{j}$, i.e., $\phi_{\mathcal{A}_{i\rightarrow j}}(a_{i})=a_{j}$, if and only if:
@@
Figure 7: H-bond path configurations for the Prism (vibration mode 00), according to their unstable sets (white: saddles, red: minima, black curves: unstable sets). A 2D slice of the electron density is shown in the background with three isolines. The corresponding partial isomorphisms are shown with colors in the insets. (a) At equilibrium ($\alpha_{10}$), the unstable set of the saddle $A$ connects it to hydrogen (left) and oxygen (right) atoms. (b) Under intermediate displacement ($\alpha_{2}$), the saddle $A$ increases its $\rho^{\prime}$ value (white isoline). The geometry of its unstable set is altered such that it now connects it to two oxygen atoms. We refer to this configuration as a misconnected H-bond. (c) Under large displacement ($\alpha_{0}$), the saddle $A$ disappears. The electron density field has been sufficiently modified by the vibrational displacement to cancel the corresponding saddle-saddle persistence pair [13]. We refer to this configuration as a missing H-bond.
@@ 5 Results
This section presents experimental results obtained with our approach, implemented in C++ within TTK [79, 7]. Our experimental data Sec. 2 is available at this address: https://github.com/thom-dani/BondMatcher. In both case studies (Secs. 5.2 and 5.3), our bond occurrence rate (Sec. 4.4) played an instrumental visual role as it enabled, with only one visualization (at equilibrium), the quick identification of the H-bonds involved in transformations (instead of exploring one by one the members of the considered ensemble). This quick identification eased their selection for refined inspections, detailed in Secs. 5.2 and 5.3.
@@ 5.1 Geometrical interpretation
Running the topological data analysis pipeline described in Sec. 3 yields a database of extremum graphs. Fig. 7 presents a classification of the H-bond configurations that we have observed in this database, according to their associated unstable sets (Sec. 1.3). This figure shows that, under vibrational displacement, the bond critical point (Sec. 1.2) associated to an H-bond path gradually increases its $\rho^{\prime}$ value. Depending on the amplitude of the vibrational displacement, this can result in two degenerate configurations. First, a misconnected H-bond, Fig. 7(b), occurs when the $\rho^{\prime}$ value of the bond critical point increased to an intermediate level, such that the geometry of its unstable set is altered to the point that it connects two oxygen atoms together (a configuration which is not chemically relevant in this context). Second, a missing H-bond, Fig. 7(c), occurs when the $\rho^{\prime}$ value of the bond critical point increased so much that the corresponding saddle-saddle pair has been canceled [13]. This figure suggests that, depending on the geometric transformation induced by the vibration, the transition from a valid H-bond path at equilibrium to a missing one is gradual (via a misconnection step), which may relate to experimental studies discussing the role of oxygen-oxygen contractions on vibrations of liquid water [84].
Fig. 6 confirms this observation by inspecting the entire set of $21$ vibrational displacements for the fist vibrational mode of the Prism. Specifically, it shows that missing H-bonds occur at the extremities of the displacement spectrum ($\alpha_{0}$ and $[\alpha_{18},\alpha_{20}]$) and that misconnected H-bonds occur within immediately adjacent intervals ($[\alpha_{1},\alpha_{3}]$ and $[\alpha_{13},\alpha_{17}]$). This confirms that the transition from valid (around the equilibrium), to misconnected, to missing H-bond is indeed gradual along this vibration. Fig. 6(c) also shows the bond occurrence rate (Sec. 4.4) for these configurations. It reports lower rates for misconnected H-bonds, indicating the transitional aspect of this configuration.
@@
Fig. 8 investigates the relation between the unstable bonds of the Prism (relative to vibrations) and topological persistence [20], a generic importance measure often used to estimate prominence and stability in Topological Data Analysis [19]. In particular, this figure shows that, in this example, the bonds identified as unstable are defined relatively to saddles which are also involved in the least persistent saddle-saddle pairs (similar observations, not reported here, were made for the Cage isomer). This indicates that such low-persistence saddles are more likely to undergo, under vibrations, the transitions observed in Fig. 7. Still, note that in theory, instability can still be observed with unstable sets defined relative to the most persistent saddles, as illustrated in Fig. 4. Then, although persistence and bond stability appear to be correlated in our experiments, it is not necessarily the case in generality.
@@
Figure 8: Observing the persistence [20] and bond occurrence rate on the Prism (equilibrium). (a) Saddle persistence pairs (persistence: black to red) : $1$-saddles create cycles in the extremum graph, while $2$-saddles represent the topological discs which fill these cycles. (b-c) Bond occurrence rate (black to red), for mode 00 (b) and mode $2$ (c). In this example, the saddles defining unstable H-bonds (arrows) are also involved in the least persistent saddle pairs (black, gray, pink).